OpenStudy (anonymous):

Determine an equation of the line of intersection of the planes 5x+y+2z+15=0 and 2x+y+2z+8=0. Answers is equation = (-7,20,0)+t(0,-2,1)

OpenStudy (anonymous):

I know how to find the direction vector (0,-2,1) but what about the point?

OpenStudy (amistre64):

5x+y+2z+15 = 2x+y+2z+8 5x +y +2z +15 -2x -y -2z -8 ----------------- 3x + 7 = 0

OpenStudy (amistre64):

x = -7/3 seems to be the line of intersection if i did it right

OpenStudy (amistre64):

i dont think I did it right :)

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

yeah i figured it out

OpenStudy (anonymous):

First we find the direction vector by crossing the normals of each line

OpenStudy (amistre64):

sounds good

OpenStudy (anonymous):

then we can find a point that is on both lines by finding a point that satisfies both equation

OpenStudy (anonymous):

im sure you probably knew all this :P

OpenStudy (anonymous):

medal still for you pops :D

OpenStudy (amistre64):

its rattling around up there someplace inbetween the rafters lol

OpenStudy (anonymous):

lol .. highschool ?

OpenStudy (amistre64):

nah..... high school went and gone many many years ago

OpenStudy (anonymous):

lol .. PhD?

OpenStudy (anonymous):

:P